OpenStudy (anonymous):

the average weight of six of seven players in the offensive line of the boston bombers is 230 pounds. what should the weight of the seventh player be to make the eintire line 239 pounds. give answer in pounds

OpenStudy (anonymous):

Average weight is just total weight divided by then number of players.

OpenStudy (anonymous):

T / 6 = 230 The new total is going to be T + x, where x is the seventh player's weight. We want (T + x) / 7 = 239

OpenStudy (anonymous):

First solve for T, then solve for x.

OpenStudy (anonymous):

how to solve for t? and x

OpenStudy (anonymous):

well solve for t by using the equation t/6 = 230 The substitute the value of t into the equation (t+x)/7 = 239

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thanks!